An urn has 12 red balls and 8 green balls. Pick 5, with replacement. What is the probability of P(3R,2B)?

-I understand how to do the problem however I do not get how we can figure out the total # of events is going to be 20^5
Answer #1

Total number of balls = 12 +8 = 20 PI 2118 CN 0100 P(G) = using binomial theorem P(3R, 26) = (?) (2) (3) - 1199920 = 0.3456